# HyDeFuse: Provably Convergent Denoiser-Driven Hyperspectral Fusion

### v1

@@ 6 Conclusion
The objective of this work was to propose an iterative denoising-based algorithm for hyperspectral fusion––modeled on the plug-and-play framework––that not only exhibits strong empirical performance but also comes with a provable convergence guarantee. Specifically, we utilized kernel denoisers, which provide effective noise reduction while maintaining tractable mathematical properties. We showed that there is scope for innovation within this class and came up with a cascaded denoiser, CasKD, which outperforms all previously reported kernel denoisers for hyperspectral images. Moreover, we demonstrated that the performance of CasKD is comparable with existing state-of-the-art (non-trained) denoisers. As explained in the introduction, we do not consider trained deep denoisers since establishing predictable mathematical properties and coming up with convergence guarantees is fundamentally challenging.
To integrate the implicit CasKD regularizer with a model-based loss for hyperspectral fusion, we employed the classical proximal gradient descent framework. Notably, leveraging the mathematical properties of CasKD, we rigorously and unconditionally established global linear convergence of our iterative fusion algorithm, HyDeFuse. We empirically validated the theoretical guarantee and analyzed how various parameters of HyDeFuse influence the convergence rate. While iterate convergence ensures the stability and reliability of the algorithm, it does not inherently guarantee high-quality reconstruction. We conducted extensive experiments on multiple hyperspectral datasets, demonstrating that HyDeFuse outperforms classical methods and performs competitively with recent state-of-the-art approaches. In particular, we observed that for proper settings of the internal parameters, the PSNR of the reconstruction stabilizes within 10–15 iterations. Although this was not discussed in the present work, the algorithmic concepts and mathematical guarantees can be extended to other models for denoising-driven regularization. It remains to be seen whether they can match state-of-the-art methods.
@@ 7.1 Proof of Proposition 4
We recall that $\mathcal{V}$ is the bandwise denoiser given by (15) and $\mathcal{W}$ is the high-dimensional denoiser given by (18). The difference is that a difference $\mathbf{W}^{(b)}$ is used for each band in $\mathcal{V}$, while the same $\mathbf{W}$ is used across all bands in $\mathcal{W}$. To show that $\mathcal{V}$ is self-adjoint, we need to check that, for all $\mathbf{X},\mathbf{Y}\in\mathbb{H}$,
$\big{\langle}\mathcal{V}(\mathbf{X}),\mathbf{Y}\big{\rangle}_{\mathbb{H}}=\big{\langle}\mathbf{X},\mathcal{V}(\mathbf{Y})\big{\rangle}_{\mathbb{H}}.$
Recall that $\mathbb{H}$ is the space of matrices of size $N_{m}\times L_{s}$. In this part, it will be convenient to use $\boldsymbol{x}_{b}$ and $\boldsymbol{y}_{b}$ for the $b$-th columns (vectors in $\mathbb{R}^{N_{m}}$) of $\mathbf{X}$ and $\mathbf{Y}$. Using this notation, we can write (4) as
@@
The verification for $\mathcal{W}$ is identical. Thus, $\mathcal{V}$ and $\mathcal{W}$ are self-adjoint.
That $\mathcal{V}$ and $\mathcal{W}$ are nonexpansive follows directly from the nonexpansive property of $\mathbf{W}^{(b)}$ and $\mathbf{W}$. Indeed, since these matrices are stochastic and symmetric, it is not difficult to show that they are nonexpansive. Hence, for any $\mathbf{X}\in\mathbb{H}$, we have
@@
where $\|\cdot\|_{2}$ is the Euclidean norm on $\mathbb{R}^{N_{m}}$. The proof for $\mathcal{W}$ is identical.
We know from the spectral theory of linear operators that a self-adjoint operator has real eigenvalues [67]. Moreover, since we have shown that $\mathcal{W}$ is nonexpansive, we can conclude that if $\lambda\in\sigma(\mathcal{W})$, then $|\lambda|\leqslant 1$. In other words, $\sigma(\mathcal{W})\subset[-1,1]$. We know from Proposition 1 that $\mathbf{W}$ is symmetric positive semidefinite. Therefore, for any $\mathbf{X}\in\mathbb{H}$,
@@
This means that every eigenvalue of $\mathcal{W}$ is nonnegative. Hence $\sigma(\mathcal{W})\subset[0,1]$.
Finally, we will show that if $\mathcal{W}(\mathbf{X})=\mathbf{X}$, then each band $\boldsymbol{x}_{b}$ must be a multiple of $\boldsymbol{e}$, the all-ones vector in $\mathbb{R}^{N_{m}}$. This follows directly from the fact that $1$ is a simple eigenvalue of $\mathbf{W}$. Indeed, since $\mathbf{W}$ is stochastic and irreducible (Proposition 1), the Perron-Frobenius theorem tells us that $1$ must be a simple eigenvalue of $\mathbf{W}$ [67]. Now, note that the equation $\mathcal{W}(\mathbf{X})=\mathbf{X}$ translates to $\mathbf{W}\boldsymbol{x}_{b}=\boldsymbol{x}_{b}$ for all $b$, i.e., $\boldsymbol{x}_{b}=\mathbf{0}$ or $\boldsymbol{x}_{b}$ is an eigenvector with eigenvalue $1$. In the latter case, $\boldsymbol{x}_{b}=c_{b}\boldsymbol{e}$ for some $c_{b}\in\mathbb{R}$. This establishes (36), also completing the proof of Proposition 4.
@@ 7.2 Proof of Lemma 1
Since $\mathcal{T}$ is a self-adjoint operator on the inner product space $\mathbb{H}$, we can diagonalize it on an orthonormal basis [67]. More specifically, if $d$ is the dimension of $\mathbb{H}$, then we can find $\lambda_{1},\ldots,\lambda_{d}\in\mathbb{R}$ and an orthonormal basis $\mathbf{H}_{1},\ldots,\mathbf{H}_{d}$ of $\mathbb{H}$ such that, for any $\mathbf{X}\in\mathbb{H}$,
@@
Let $\mathbf{X}\in\mathbb{H}$. We have from (39) that $\|\mathcal{T}(\mathbf{X})\|_{\mathbb{H}}=\|\mathbf{X}\|_{\mathbb{H}}$ if and only if
@@
However, $\lambda_{j}\in(-1,1]$ by assumption. Therefore, we must have $\lambda_{j}=1$ or $\langle\mathbf{X},\mathbf{H}_{j}\rangle_{\mathbb{H}}=0$. In either case, it follows from (38) that $\mathcal{T}(\mathbf{X})=\mathbf{X}$. This completes the proof of Lemma 1.

### v2

@@ 6 Conclusion
We developed an iterative, denoising-based algorithm for hyperspectral fusion within the plug-and-play (PnP) framework. The proposed method combines strong empirical performance with provable convergence guarantees. This is achieved through the use of kernel denoisers, which provide effective noise suppression while maintaining mathematically tractable properties. In particular, we introduced CasKD, a cascaded kernel denoiser that captures both inter-band and intra-band correlations in hyperspectral data. CasKD outperforms previously reported kernel denoisers and performs on par with leading classical (non-trained) methods. We did not consider learned deep denoisers, as ensuring predictable mathematical behavior and convergence guarantees for such models remains fundamentally challenging.
To integrate the implicit CasKD regularizer with a model-based loss for hyperspectral fusion, we employed the classical proximal gradient descent (PGD) framework. We then established a rigorous operator-theoretic formulation of CasKD and the resulting fusion algorithm, HyDeFuse, providing a clear connection between the denoiser properties and the convergence behavior of the iterative scheme. Based on this formulation, we derived an unconditional guarantee of global linear convergence for HyDeFuse.
While this result ensures algorithmic stability and reliability, it does not by itself guarantee high-quality reconstructions. To validate the practical performance, we conducted extensive experiments on multiple hyperspectral datasets, showing that HyDeFuse consistently outperforms classical approaches and remains competitive with recent state-of-the-art methods. With suitable parameter settings, the PSNR stabilizes within 10–15 iterations. The underlying principles and convergence analysis can naturally extend to other denoising-based regularization models.
Data Availability Statement
All data that support the findings of this study are included within the article. The code used in this article will be made publicly available at https://github.com/sagarkr1729/Contractive-Hyperspectral-Fusion upon acceptance of this manuscript.
@@ 7 Appendix
In this section, we prove 3, 4, 5, 6, and Lemma 1.
7.1 Proof of 3
We first show that $\mathcal{K}$ is self-adjoint on $\mathbb{H}$. Indeed, for any $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{H}$, we have
$\big\langle\mathcal{K}(\mathbf{X}_{1}),\mathbf{X}_{2}\big\rangle_{\mathbb{H}}$ $=$ $\mathrm{tr}\big((\mathbf{P}_{1}\mathbf{X}_{1}+\mathbf{X}_{1}\mathbf{P}_{2})^{\top}\mathbf{X}_{2}\big)$
where the cyclic property of trace and the symmetry of $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ are used. Hence, $\mathcal{K}$ is self-adjoint.
We next show that $\mathcal{K}$ is $\beta$-Lipschitz, where
$\beta=\sigma_{\max}(\mathbf{A})^{2}+\lambda\,\sigma_{\max}(\mathbf{E}\mathbf{R})^{2}.$ (31)
From (12) and the triangle inequality,
$\|\mathcal{K}(\mathbf{X})\|_{\mathbb{H}}\leqslant\|\mathbf{P}_{1}\mathbf{X}\|_{\mathbb{H}}+\|\mathbf{X}\mathbf{P}_{2}\|_{\mathbb{H}}.$
Let $\mathbf{P}_{1}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}$ be the SVD of $\mathbf{P}_{1}$, where $\mathbf{U}$ and $\mathbf{V}$ are orthogonal, and $\mathbf{\Sigma}=\mathrm{diag}(\sigma_{1},\dots,\sigma_{r})$ are the singular values. Since orthogonal matrices preserve the Frobenius norm,
$\|\mathbf{P}_{1}\mathbf{X}\|_{\mathbb{H}}=\|\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}\mathbf{X}\|_{\mathbb{H}}=\|\mathbf{\Sigma}\mathbf{Y}\|_{\mathbb{H}},\qquad\mathbf{Y}:=\mathbf{V}^{\top}\mathbf{X}.$
Because $\mathbf{\Sigma}$ is diagonal,
$\|\mathbf{\Sigma}\mathbf{Y}\|_{\mathbb{H}}^{2}=\sum_{i,j}\sigma_{i}^{2}|\mathbf{Y}_{ij}|^{2}\leqslant\sigma_{\max}(\mathbf{P}_{1})^{2}\sum_{i,j}|\mathbf{Y}_{ij}|^{2}=\sigma_{\max}(\mathbf{P}_{1})^{2}\|\mathbf{Y}\|_{\mathbb{H}}^{2}.$
Since $\mathbf{V}$ is orthogonal, $\|\mathbf{Y}\|_{\mathbb{H}}=\|\mathbf{X}\|_{\mathbb{H}}$, giving
$\|\mathbf{P}_{1}\mathbf{X}\|_{\mathbb{H}}\leqslant\sigma_{\max}(\mathbf{P}_{1})\|\mathbf{X}\|_{\mathbb{H}}.$
Similarly, taking transposes,
$\|\mathbf{X}\mathbf{P}_{2}\|_{\mathbb{H}}=\|\mathbf{P}_{2}^{\top}\mathbf{X}^{\top}\|_{\mathbb{H}}\leqslant\sigma_{\max}(\mathbf{P}_{2})\|\mathbf{X}^{\top}\|_{\mathbb{H}}=\sigma_{\max}(\mathbf{P}_{2})\|\mathbf{X}\|_{\mathbb{H}}.$
Combining these bounds, we obtain
$\|\mathcal{K}(\mathbf{X})\|_{\mathbb{H}}\leqslant\beta\|\mathbf{X}\|_{\mathbb{H}},\quad\beta:=\sigma_{\max}(\mathbf{P}_{1})+\sigma_{\max}(\mathbf{P}_{2}).$ (32)
Hence, $\mathcal{K}$ is $\beta$-Lipschitz. Substituting $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ from (12) yields the expression for $\beta$ in (31).
@@ 7.2 Proof of Proposition 4
From (11), we have
Hence, by (32), $\nabla\!\ell$ is $\beta$-smooth. Moreover, since $\ell$ is convex, the Baillon–Haddad theorem [74] gives
$\langle\nabla\!\ell(\mathbf{X}_{1})-\nabla\!\ell(\mathbf{X}_{2}),\mathbf{X}_{1}-\mathbf{X}_{2}\ \rangle_{\mathbb{H}}\geqslant\frac{1}{\beta}\|\nabla\!\ell(\mathbf{X}_{1})-\nabla\!\ell(\mathbf{X}_{2})\|_{\mathbb{H}}^{2}\qquad(\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{H}).$ (34)
Substituting (33), we obtain
From (26), we can write
$\|\mathcal{G}(\mathbf{X})\|_{\mathbb{H}}^{2}=\|\mathbf{X}\|_{\mathbb{H}}^{2}+\gamma^{2}\|\mathcal{K}(\mathbf{X})\|_{\mathbb{H}}^{2}-2\gamma\langle\mathbf{X},\mathcal{K}(\mathbf{X})\ \rangle_{\mathbb{H}}.$
Substituting (35) gives
Hence, if $0<\gamma\leqslant 2/\beta$, we have $\|\mathcal{G}(\mathbf{X})\|_{\mathbb{H}}\leqslant\|\mathbf{X}\|_{\mathbb{H}}$. Since $\mathcal{G}\in\mathcal{L}(\mathbb{H})$, it follows that $\mathcal{G}$ is nonexpansive.
To complete the proof of Proposition (4), we must show that 1-1 is not an eigenvalue of $\mathcal{G}$. Indeed, if $1\in\sigma(\mathcal{G})$, then there exists a nonzero $\overline{\mathbf{X}}\in\mathbb{H}$ such that
However, this implies $2\overline{\mathbf{X}}=\gamma\mathcal{K}(\overline{\mathbf{X}})$, so that if $0<\gamma<2/\beta$, we have from (32) that
which is a contradiction. Therefore, we must have $1\notin\sigma(\mathcal{G})$.
7.3 Proof of Proposition 5
Recall that $\mathcal{V}$ is the bandwise denoiser defined in (16), while $\mathcal{W}$ is the high-dimensional denoiser defined in (19). The key distinction is that $\mathcal{V}$ applies a separate denoiser $\mathbf{W}^{(b)}$ to each spectral band (see (15)), whereas $\mathcal{W}$ uses a single shared denoiser $\mathbf{W}$ across all bands (see (18)).
To show that $\mathcal{V}$ is self-adjoint, we need to verify that
$\big\langle\mathcal{V}(\mathbf{X}),\mathbf{Y}\big\rangle_{\mathbb{H}}=\big\langle\mathbf{X},\mathcal{V}(\mathbf{Y})\big\rangle_{\mathbb{H}}\qquad(\mathbf{X},\mathbf{Y}\in\mathbb{H}).$
Recall that $\mathbb{H}$ is the space of matrices of size $N_{m}\times L_{s}$. It will be convenient to denote by $\boldsymbol{x}_{b}$ and $\boldsymbol{y}_{b}$ the $b$-th columns (vectors in $\mathbb{R}^{N_{m}}$) of $\mathbf{X}$ and $\mathbf{Y}$, respectively. With this notation, we can rewrite (4) as
@@ 7.3 Proof of Proposition 5
The verification for $\mathcal{W}$ is identical.
Since $\mathbf{W}^{(b)}$ and $\mathbf{W}$ are stochastic and symmetric, it is straightforward to verify that they are nonexpansive. Consequently, the associated operators $\mathcal{V}$ and $\mathcal{W}$ are also nonexpansive. Indeed, for any $\mathbf{X}\in\mathbb{H}$, we have
@@
where $\|\cdot\|_{2}$ is the $\ell_{2}$ norm on $\mathbb{R}^{N_{m}}$. The argument for $\mathcal{W}$ is identical.
We know that a self-adjoint operator has real eigenvalues [70]. Moreover, since $\mathcal{W}$ has been shown to be nonexpansive, it follows that if $\lambda\in\sigma(\mathcal{W})$, then $|\lambda|\leqslant 1$, i.e., $\sigma(\mathcal{W})\subset[-1,1]$. Further, from Proposition 1, $\mathbf{W}$ is symmetric and positive semidefinite. Hence, for any $\mathbf{X}\in\mathbb{H}$,
@@
This shows that all eigenvalues of $\mathcal{W}$ are nonnegative, and therefore $\sigma(\mathcal{W})\subset[0,1]$.
Finally, we show that if $\mathcal{W}(\mathbf{X})=\mathbf{X}$, then each band $\boldsymbol{x}_{b}$ must be a multiple of $\boldsymbol{e}$, the all-ones vector in $\mathbb{R}^{N_{m}}$. This follows directly from the fact that $1$ is a simple eigenvalue of $\mathbf{W}$. Indeed, the equation $\mathcal{W}(\mathbf{X})=\mathbf{X}$ translates to $\mathbf{W}\boldsymbol{x}_{b}=\boldsymbol{x}_{b}$ for all $b$, i.e., $\boldsymbol{x}_{b}$ is an eigenvector with eigenvalue $1$. Since $\mathbf{W}$ is stochastic and irreducible (Proposition 1), the Perron-Frobenius theorem tells us that $1$ must be a simple eigenvalue of $\mathbf{W}$ [70]. Consequently, we must have $\boldsymbol{x}_{b}=c_{b}\boldsymbol{e}$ for some $c_{b}\in\mathbb{R}$. This establishes (27) and completes the proof of Proposition 5.
7.4 Proof of Proposition 6
From Proposition 5, we know that $\mathrm{fix}\,(\mathcal{W})$ consists of matrices $\mathbf{F}$ of the form (27). Therefore, to establish the stated result, it suffices to show that if $\mathbf{F}$ is of the form (27) and $\mathbf{F}\neq\mathbf{0}$, then $\mathbf{F}\notin\mathrm{fix}\,(\mathcal{G})$.
We will show, by contradiction, that under this condition $\mathbf{F}\notin\mathrm{fix}\,(\mathcal{G})$. Suppose that $\mathcal{G}(\mathbf{F})=\mathbf{F}$. Then, from (26), we have $\mathcal{K}(\mathbf{F})=\mathbf{0}$, i.e.,
$\mathbf{P}_{1}\mathbf{F}+\mathbf{F}\mathbf{P}_{2}=\mathbf{0},$
Since $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ are positive semidefinite, we know that the two terms on the right are nonnegative. Therefore, we must have
where $\|\cdot\|_{2}$ is the $\ell_{2}$ norm on $\mathbb{R}^{N_{m}}$. Now, since $\mathbf{F}\neq\mathbf{0}$, we have $c_{j}\neq 0$ for some $1\leqslant j\leqslant L_{s}$. Hence, it follows that $\mathbf{A}\boldsymbol{e}=\mathbf{0}$.
Recall that $\mathbf{A}=\mathbf{S}\mathbf{B}$, where $\mathbf{S}$ and $\mathbf{B}$ are sampling and blur operators. Since $\mathbf{B}\boldsymbol{e}=\boldsymbol{e}$, i.e., blurring a constant-intensity image produces the same image, we obtain $\mathbf{0}=\mathbf{A}\boldsymbol{e}=\mathbf{S}(\mathbf{B}\boldsymbol{e})=\mathbf{S}\boldsymbol{e}$. This leads to a contradiction, as sampling $\boldsymbol{e}$ cannot yield the zero vector.
@@ 7.5 Proof of Lemma 1
Since $\mathcal{T}$ is a self-adjoint operator on the inner product space $\mathbb{H}$, it can be diagonalized with respect to an orthonormal basis [70]. Specifically, if $d$ is the dimension of $\mathbb{H}$, then we can find $\lambda_{1},\ldots,\lambda_{d}\in\mathbb{R}$ and an orthonormal basis $\mathbf{H}_{1},\ldots,\mathbf{H}_{d}$ of $\mathbb{H}$ such that, for any $\mathbf{X}\in\mathbb{H}$,
@@
Now, from (38), we have that $\|\mathcal{T}(\mathbf{X})\|_{\mathbb{H}}=\|\mathbf{X}\|_{\mathbb{H}}$ if and only if
@@
Since $\lambda_{j}\in(-1,1]$ by assumption, it follows that either $\lambda_{j}=1$ or $\langle\mathbf{X},\mathbf{H}_{j}\rangle_{\mathbb{H}}=0$. In both cases, it follows from (37) that $\mathcal{T}(\mathbf{X})=\mathbf{X}$. This completes the proof of Lemma 1.